OpenStudy (anonymous):

MEDAL & FAN ANYBODY ? If the height of a cone is tripled and radius is held constant, then the volume will double. True False

OpenStudy (anonymous):

@sleepyjess

OpenStudy (hauntedwoodsgal):

True

OpenStudy (sleepyjess):

lets say our height is 9 and radius is 5 for the first cone which means the height is 27 and radius is 5 for the second cone \(V=\pi r^2\dfrac{h}{3}\)

OpenStudy (sleepyjess):

\(\pi 5^2\dfrac{9}{3} \\5^2=25\\\dfrac{9}{3}=3\\3*25=75\\75*\pi=235.5\)

OpenStudy (anonymous):

thank you :)

OpenStudy (sleepyjess):

no problem \(\huge\ddot\smile\)

OpenStudy (sleepyjess):

how did you come up with your answer @Hauntedwoodsgal